I need help with this word problem please; find the distance from third base to first base on a baseball diamond if the distance between bases is 90 feet.
Answer by checkley75(3666)

WE HAVE A RIGHT TRIANGLE WITH SIDES OF 90 FEET & ARE LOOKING FOR THE HYPOTENUSE.
A^2+B^2=C^2
90^2+90^2=C^2
8100+8100=C^2
16200=C^2
C=SQRT16200
C=127.279 ANSWER.